2. Write each of the following quadratic functions in the form y=ax^2 + bx + c.
a) y= 2x^2 - 7
b) f(x) = 4(x-1)(x+3)
=4x^2 + 2x - 3(Is this CORRECT???)
c) 2(y-4)+7x=-3(x+1)^2-5
Kinda unsure of this whole question. Thanks! Help is appreciated.

Answer by htmentor(1343)

(a) is already in the correct form.
For (b), multiply the two product terms using FOIL, then multiply through by 4:
f(x) = 4(x^2 + 2x - 3) = 4x^2 + 8x - 12
[Your answer would have been correct, except you didn't multiply all the terms by '4']
For (c), carry out the multiplication and collect like terms:
2y - 8 + 7x = -3(x^2 + 2x + 1) - 5
2y - 8 + 7x = -3x^2 - 6x - 3 - 5
2y = -3x^2 - 13x - 8 + 8
So we have y = (-3/2)x^2 - (13/2)x
Let's check our answer to (c) with a simple example. Let x = 1
Original equation gives 2(y-4) + 7 = -3(2)^2 - 5 -> 2y - 8 + 7 = -12 - 5 -> y = -8
Solved equation gives y = (-3/2)(1)^2 - (13/2)(1) - > y = -16/2 = -8